8 Compa hy be ught and staxt to use. Ci Sse escet asset was coshnIsC0o,with an esitimateals 5 oc o hours?// Yeg a 1..-3 coe ho.uXS/.year 7ooo, Compare the results a the n the hles thion yemethod. I thEll, pre ciaboh- method?.hy Ell i Sthaighu line methed preciaboh Do Depreciation

please solve this in straight line method,sum of the year nd performance related method
Answer #1
Straight Line method of Depreciation in $
Cost of the asset 150000
Salvage Value 30000
Estimated life of the asset 2 years
Depreciation per annum = Cost of the asset (-) Scrap Value
Estimated life of the asset
= 150000-30000
2
= 120000/2
= 60000 per annum
Year Depreciation Accumulated Deprecaition Net Book Value
1                                       60,000 60000 150000-60000 90000
2 60000 120000 150000-120000 30000 Residual Value
Sum of Years Digits Method in $
Cost of the asset 150000
Salvage Value 30000
Estimated life of the asset 2 years
Sum of Years Digits (SYD) = n(n+1)/2
Where n = estimated useful life
SYD = 2(2+1) / 2
= 3
Year Balance of useful life at the begininning of the year SYD % Annual Depreciation
1 2 2/3 66.67% 80004 (120000 x 66.67%)
2 1 1/3 33.33% 39996 (120000 x 33.33%)
100.00% 120000 (150000-30000 Residual value)
Performance method of Depreciation in $
Cost of the asset 150000
Salvage Value 30000
Estimated life of the asset 15000 Hours
(In Hours)
Depreciation per hour = Cost of asset (-) Scrap Value
No of Estimated life in hours
= 150000 (-) 30000
15000 Hours
= 120000 /15000
Per hour rate of Depreciation= 8
Year Depreciation Accumulated Deprecaition Net Book Value
(No of Hours x Rate per Hour)
1                                       64,000 64000 150000-64000 86000
(8000 x8)
2 56000 120000 150000-120000 30000 Residual Value
(7000 x 8)